# Einstein metrics on homogeneous superspaces

### v1

@@ 1 Introduction
Over the past few decades, considerable efforts have been made to extend Riemannian geometry to the realm of supermanifolds; see, for example, the classic book [DeW92] and the more recent texts [Rog07, CCF11, Kes19]. Notions such as Riemannian metric, Levi-Civita connection and Riemann curvature have thus been generalised to this setting. The purpose of the present paper is to address some of the fundamental mathematical aspects of these notions, building on explicit curvature formulas that we derive for certain types of supermanifolds. The appearance of supermanifolds in contemporary physical theories (see, e.g., [MT98, BBHZZ00, AF09, Zar10, Wit19a, Wit19b, Kes19] and references therein) makes their mathematical study timely and warrants the development of manageable curvature expressions in cases with physically relevant symmetries or properties.
@@
Another goal in Riemannian geometry is the exploration of links between the geometry and the topology of a manifold. For instance, the Bonnet–Myers theorem states that the positivity of the Ricci curvature implies the finiteness of the fundamental group. In the presence of symmetry, the links are particularly strong. A compact homogeneous space, for example, typically does not admit a metric with non-positive Ricci curvature. According to the Bochner theorem (see [Bes87, Theorem 1.84]), if $G_{0}/K_{0}$ is a compact homogeneous space with non-positive Ricci curvature, then the connected component of the identity of $G_{0}$ must be abelian.
@@
Section 4 is devoted to Einstein metrics on homogeneous superspaces $G/K$, where $G$ is a connected Lie supergroup with Lie superalgebra $\mathfrak{g}$. We assume that $\mathfrak{g}$ is basic classical to ensure that $\mathfrak{g}^{\mathbb{C}}$ admits a non-degenerate $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariant supersymmetric even bilinear form, which we use to obtain our background metric on $G/K$. Furthermore, we demand that $\mathfrak{g}$ be the compact real form of $\mathfrak{g}^{\mathbb{C}}$ induced by a so-called star operation, as is the case in the non-super setting when $G_{0}$ is compact. This leaves $\mathfrak{g}^{\mathbb{C}}$ to be of type $A$ or $C$, whence $G=\operatorname{SU}(m|n)$ or $G=\operatorname{SOSp}(2|2n)$; see [SNR77, AN15, FG23]. Imitating the structure of so-called generalised flag manifolds in classical Riemannian geometry (see [Arv03]), we will assume that the Dynkin diagram of $K$ is obtained from that of $G$ by removing one or more nodes. Here, we restrict ourselves to removing one or two nodes only but intend to discuss elsewhere the removal of an arbitrary number of nodes. Section 4.1 thus outlines the construction of homogeneous superspaces by means of Dynkin diagrams, while Section 4.2 concerns the corresponding flag supermanifolds and their compact real forms. In Section 4.3 and Section 4.4, we obtain our classification results for Einstein metrics on $G/K$ with $G=\operatorname{SU}(m|n)$. Notably, some of the metrics we find are Ricci-flat. Section 4.5 concerns Einstein metrics on $G/K$ with $G=\operatorname{SOSp}(2|2n)$.
@@
Throughout, the ground field is the field $\mathbb{R}$ of real numbers, unless otherwise stated. We let $\mathbb{C}$ denote the field of complex numbers, $\mathbb{H}$ the quaternions, $\mathbb{Z}$ the ring of integers, $\mathbb{R}^{+}$ (respectively $\mathbb{R}^{-}$) the set of positive (respectively negative) real numbers, and $\imath$ the imaginary unit, and set ×=$\mathbb{R}^{\times}=\mathbb{R}-\{0\}$. If $\{v_{1},\ldots,v_{d}\}$ is a basis for a vector superspace with scalar superproduct $\langle$-,$\rangle$ - , - ⟩, then we let $\{\bar{v}_{1},\ldots,\bar{v}_{d}\}$ denote the corresponding right dual basis, i.e., $\langle v_{i},\bar{v}_{j}\rangle=\delta_{ij}$ for all $i,j\in\{1,\ldots,d\}$.
@@ 2 Supermanifolds
Here, we recall some basics on supergeometry and refer to [Kos77, Bat79, Lei80, BBH91, DeW92, DM99, Var04, Goe08, CCF11] for more details. Focus is on notions and results needed in subsequent sections.
@@ 2.1 Lie superalgebras and real forms
A vector superspace is a $\mathbb{Z}_{2}$-graded vector space $V=V_{\bar{0}}\oplus V_{\bar{1}}$, where $\mathbb{Z}_{2}=\{\bar{0},\bar{1}\}$. The superdimension of $V$ is defined as $\operatorname{sdim}(V)=m-n$, where $m=\dim(V_{\bar{0}})$ and $n=\dim(V_{\bar{1}})$. A nonzero element of $V_{\bar{0}}\cup V_{\bar{1}}$ is said to be homogeneous, with elements of $V_{\bar{0}}$ called even and those of $V_{\bar{1}}$ odd, and a basis $\{v_{1},\dots,v_{m+n}\}$ for $V$ is said to be homogeneous if every $v_{i}$ is homogeneous. A homogeneous vector $v\in V_{a}$ is said to have parity $[v]=a$. For the tensor product of the vector superspaces $V$ and $W$, we have
@@
Throughout this paper, definitions and calculations for vector superspaces are given in terms of homogeneous elements and then extended linearly. Moreover, the ground field $\mathbb{R}$ is considered an even vector space over $\mathbb{R}$, and a linear map $f:V\to W$ is said to be even if $f(V_{a})\subseteq W_{a}$ for $a\in\mathbb{Z}_{2}$. This notion extends readily to bilinear maps.
A Lie superalgebra is a $\mathbb{Z}_{2}$-graded vector space (over $\mathbb{R}$ or $\mathbb{C}$) equipped with a Lie superbracket [,][-,-][ - , - ] which is super-skew-symmetric and satisfies the super-Jacobi identity. The (even bilinear) Killing form is defined for every finite-dimensional Lie superalgebra $\mathfrak{g}$, by
@@
A finite-dimensional complex Lie superalgebra $\mathfrak{g}^{\mathbb{C}}=\mathfrak{g}_{\bar{0}}^{\mathbb{C}}\oplus\mathfrak{g}_{\bar{1}}^{\mathbb{C}}$ is called basic classical if $\mathfrak{g}^{\mathbb{C}}$ is simple, $\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$ is reductive, and $\mathfrak{g}^{\mathbb{C}}$ admits a non-degenerate $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariant supersymmetric even bilinear form. Following Kac [Kac77, Kac78], there are two types of basic classical Lie superalgebras, characterised by whether $\mathfrak{g}_{\bar{1}}^{\mathbb{C}}$ is reducible (type \@slowromancapi@) or irreducible (type \@slowromancapii)@ as a $\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$-module. Thus, $A(m,n)$ and $C(n)$ comprise the basic classical Lie superalgebras of type \@slowromancapi@, while $B(m,n)$, $D(m,n)$, $F(4)$, $G(3)$ and $D(2,1,\alpha)$ comprise the ones of type \@slowromancapii@. Concretely, a basic classical Lie superalgebra of type \@slowromancapi@ is isomorphic to one of the following Lie superalgebras: $A(m,n)=\mathfrak{sl}(m+1|n+1)^{\mathbb{C}}$ with $m>n\geq 0$; $A(n|n)=\mathfrak{psl}(n+1|n+1)^{\mathbb{C}}$ with $n\geq 1$; or $C(n)=\mathfrak{osp}(2|2n-2)^{\mathbb{C}}$ with $n\geq 3$ (noting that $C(2)\cong A(1|0)$).
A real Lie superalgebra $\mathfrak{g}$ is called a real form of a complex Lie superalgebra $\mathfrak{g}^{\mathbb{C}}$ if $\mathfrak{g}^{\mathbb{C}}\cong\mathfrak{g}\otimes_{\mathbb{R}}\mathbb{C}$, while a star operation $*:\mathfrak{g}^{\mathbb{C}}\to\mathfrak{g}^{\mathbb{C}}$ is an even anti-linear map such that
@@
for any $X,Y\in\mathfrak{g}^{\mathbb{C}}$; see [SNR77] or [DM99, §4.1–4.3]. Here, anti-linearity means that $(cX)^{*}=\bar{c}X^{*}$, where $\bar{c}$ is its complex conjugate of $c\in\mathbb{C}$. With $\sqrt{\imath}=e^{\frac{\pi}{4}\imath}$, a real form $\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}}$ can now be constructed as
$\begin{gathered}\mathfrak{g}_{\bar{0}}=\operatorname{span}_{\mathbb{R}}\{X\in\mathfrak{g}_{\bar{0}}^{\mathbb{C}}\mid X^{*}=-X\},\\[2.84544pt] \mathfrak{g}_{\bar{1}}=\operatorname{span}_{\mathbb{R}}\{\sqrt{\imath}X\in\mathfrak{g}_{\bar{1}}^{\mathbb{C}}\mid X^{*}=-X\}=\operatorname{span}_{\mathbb{R}}\{X\in\mathfrak{g}_{\bar{1}}^{\mathbb{C}}\mid X^{*}=\imath X\}.\end{gathered}$ (2.2)
Note that $\mathfrak{g}_{\bar{0}}$ is a real form of the complex Lie algebra $\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$. We say that a real form of $\mathfrak{g}^{\mathbb{C}}$ is compact if $\mathfrak{g}_{\bar{0}}$ is a compact Lie algebra, cf. [AN15, FG23]. It was shown in [SNR77] that among basic classical Lie superalgebras, only type I Lie superalgebras admit star operations. In that case, there are actually two star operations, and they are related by duality. Throughout this paper, we are only concerned with the so-called type (1) star operation; see, e.g., [GZ90a, GZ90b] and Section 4.2.
@@ 2.2 Lie supergroups
for any open set $U\subseteq\mathbb{R}^{m}$. Here, $\mathcal{C}_{\mathbb{R}^{m}}^{\infty}$ denotes the sheaf of smooth functions on $\mathbb{R}^{m}$, and $\Lambda_{\mathbb{R}}[\xi_{1},\dots,\xi_{n}]$ is the Grassman algebra in the odd variables $\xi_{1},\dots,\xi_{n}$. A (real smooth) supermanifold of dimension $m|n$ is a ringed space $M=(|M|,\mathcal{O}_{M})$, where the underlying topological space $|M|$ is Hausdorff and second-countable, and $M$ is locally homeomorphic to $\mathbb{R}^{m|n}$, i.e., for each $x\in|M|$ there exists an open neighbourhood $V_{x}\subseteq|M|$ of $x$ and an open set $U\subseteq\mathbb{R}^{m|n}$ such that $\mathcal{O}_{M}(V_{x})\cong\mathcal{O}_{\mathbb{R}^{m|n}}(U)$. Sections of the sheaf $\mathcal{O}_{M}$ are called superfunctions on $M$, and the nilpotent such sections generate an ideal $\mathcal{I}_{M}$ in $\mathcal{O}_{M}$. The corresponding subspace $M_{0}=(|M|,\mathcal{O}_{M}/\mathcal{I}_{M})$ is called the reduced manifold associated with $M$. Note that $M_{0}$ lies in the category of ordinary smooth manifolds.
@@
$\varphi(ab)=\varphi(a)b+(-1)^{[\varphi][a]}a\varphi(b),\qquad a,b\in\mathcal{O}_{M}(U).$
The sections of $\mathcal{T}_{M}$ are called vector fields. For every point $p\in|M|$, the tangent superspace $T_{p}M$ at $p$ is defined as the vector superspace of super-derivations of the stalk $\mathcal{O}_{M,p}$, i.e., linear maps $\psi:\mathcal{O}_{M,p}\to\mathbb{R}$ such that
$\psi(fg)=\psi(f)g(p)+(-1)^{[\psi][f]}f(p)\psi(g).$
@@
$X_{e}\mapsto X=(1\otimes X_{e})\circ m^{*}.$
We refer to [CCF11] for more details.

### v2

@@ 1 Introduction
Over the past few decades, considerable efforts have been made to extend Riemannian geometry to the realm of supermanifolds; see, for example, the classic book [DeW92] and the more recent texts [Rog07, CCF11, Kes19]. Notions such as Riemannian metric, Levi-Civita connection and Riemann curvature have thus been generalised to this setting. The purpose of the present paper is to address some of the fundamental mathematical aspects of these notions, building on explicit curvature formulas that we derive for certain types of supermanifolds.
[…]
@@
Another goal in Riemannian geometry is the establishment and exploration of links between the geometry and the topology of a manifold. For instance, the Bonnet–Myers theorem states that the positivity of the Ricci curvature implies the finiteness of the fundamental group. In the presence of symmetry, the links are particularly strong. A compact homogeneous space, for example, typically does not admit a metric with non-positive Ricci curvature. According to the Bochner theorem (see [Bes87, Theorem 1.84]), if $G_{0}/K_{0}$ is a compact homogeneous space with non-positive Ricci curvature, then the connected component of the identity of $G_{0}$ must be abelian.
@@
Section 4 is devoted to Einstein metrics on homogeneous superspaces $G/K$, where $G$ is a connected Lie supergroup with Lie superalgebra $\mathfrak{g}$. We assume that $\mathfrak{g}$ is basic classical to ensure that $\mathfrak{g}^{\mathbb{C}}$ admits a non-degenerate $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariant supersymmetric even bilinear form, which we use to obtain our background metric on $G/K$. Furthermore, we demand that $\mathfrak{g}$ be the compact real form of $\mathfrak{g}^{\mathbb{C}}$ induced by a so-called star operation, as is the case in the non-super setting when $G_{0}$ is compact. This leaves $\mathfrak{g}^{\mathbb{C}}$ to be of type $A$ or $C$, whence $G=\operatorname{SU}(m|n)$ or $G=\operatorname{SOSp}(2|2n)$; see [SNR77, AN15, FG23]. Imitating the structure of so-called generalised flag manifolds in classical Riemannian geometry (see [Arv03]), we will assume that the Dynkin diagram of $K$ is obtained from that of $G$ by removing one or more nodes. Here, we restrict ourselves to removing one or two nodes only but intend to discuss elsewhere the removal of an arbitrary number of nodes. Section 4.1 thus outlines the construction of homogeneous superspaces by means of Dynkin diagrams, while Section 4.2 concerns the corresponding flag supermanifolds and their compact real forms. In Section 4.3 and Section 4.4, we obtain our classification results for Einstein metrics on $G/K$ with $G=\operatorname{SU}(m|n)$. Notably, some of the metrics we find are Ricci-flat, and our construction in Section 4.3 produces supergeometric analogues of Fubini–Study metrics on projective spaces. Section 4.5 concerns Einstein metrics on $G/K$ with $G=\operatorname{SOSp}(2|2n)$.
@@
Throughout, the ground field is the field $\mathbb{R}$ of real numbers, unless otherwise stated. We let $\mathbb{C}$ denote the field of complex numbers, $\mathbb{H}$ the quaternions (appearing in Section 4.2.2), $\mathbb{Z}$ the ring of integers, $\mathbb{Z}_{2}=\{\bar{0},\bar{1}\}$ the ring of integers modulo 2, $\mathbb{R}^{+}$ (respectively $\mathbb{R}^{-}$) the set of positive (respectively negative) real numbers, and $\imath$ the imaginary unit, and set $\mathbb{R}^{\times}=\mathbb{R}-\{0\}$. Let $\langle-,$-$\rangle$ denote a scalar superproduct on a vector superspace $V$. If $\{v_{1},\ldots,v_{d}\}$ is a basis for $V$, we write $\{\bar{v}_{1},\ldots,\bar{v}_{d}\}$ for the corresponding right dual basis, i.e., $\langle v_{i},\bar{v}_{j}\rangle=\delta_{ij}$ for all $i,j\in\{1,\ldots,d\}$. We denote by $g$ a graded Riemannian metric on a Riemannian supermanifold $M$. We write $\mathfrak{g}^{\mathbb{C}}$ for a complex Lie superalgebra, with $\mathfrak{g}$ its real form over $\mathbb{R}$. If $\mathfrak{g}^{\mathbb{C}}$ is a basic classic complex Lie superalgebra, we denote by $Q$ a fixed, non-degenerate, supersymmetric, invariant, even bilinear form on $\mathfrak{g}^{\mathbb{C}}$.
@@ 2 Supermanifolds
Here, we recall some basic notions from supergeometry. We refer to [Kos77, Bat79, Lei80, BBH91, DeW92, DM99, Var04, Goe08, CCF11] for further background, noting (as in [BCF10]) that these references employ different approaches to supergeometry. Focus here is on notions and results needed in subsequent sections.
@@ 2.1 Lie superalgebras and real forms
A vector superspace of dimension $m|n$ is a $\mathbb{Z}_{2}$-graded vector space $V=V_{\bar{0}}\oplus V_{\bar{1}}$, where $m=\dim(V_{\bar{0}})$ and $n=\dim(V_{\bar{1}})$. The superdimension of $V$ is defined as $\operatorname{sdim}(V)=m-n$. A nonzero element of $V_{\bar{0}}\cup V_{\bar{1}}$ is said to be homogeneous, with elements of $V_{\bar{0}}$ called even and those of $V_{\bar{1}}$ odd, and a basis $\{v_{1},\dots,v_{m+n}\}$ for $V$ is said to be homogeneous if every $v_{i}$ is homogeneous. A homogeneous vector $v\in V_{a}$ is said to have parity $[v]=a$. For the tensor product of vector superspaces $V$ and $W$, we have
@@
Throughout this paper, definitions and calculations for vector superspaces are given in terms of homogeneous elements and then extended linearly. Moreover, the ground field $\mathbb{R}$ is considered an even vector space over $\mathbb{R}$, and a linear map $f\in\mathrm{Hom}_{\mathbb{R}}(V,W)$ is said to be even if $f(V_{a})\subseteq W_{a}$ for $a\in\mathbb{Z}_{2}$. This notion extends readily to bilinear maps.
A Lie superalgebra is a $\mathbb{Z}_{2}$-graded vector space (over $\mathbb{R}$ or $\mathbb{C}$) equipped with a Lie superbracket [,][-,-] which is super-skew-symmetric and satisfies the super-Jacobi identity. The (even bilinear) Killing form is defined for every finite-dimensional Lie superalgebra $\mathfrak{g}$, by
@@
A finite-dimensional complex Lie superalgebra $\mathfrak{g}^{\mathbb{C}}=\mathfrak{g}_{\bar{0}}^{\mathbb{C}}\oplus\mathfrak{g}_{\bar{1}}^{\mathbb{C}}$ is called basic classical if $\mathfrak{g}^{\mathbb{C}}$ is simple, $\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$ is reductive, and $\mathfrak{g}^{\mathbb{C}}$ admits a non-degenerate $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariant supersymmetric even bilinear form. Following Kac [Kac77, Kac78], there are two types of basic classical Lie superalgebras, characterised by whether $\mathfrak{g}_{\bar{1}}^{\mathbb{C}}$ is reducible (type I) or irreducible (type II) as a $\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$-module. Thus, $A(m,n)$ and $C(n)$ comprise the basic classical Lie superalgebras of type I, while $B(m,n)$, $D(m,n)$, $F(4)$, $G(3)$ and $D(2,1,\alpha)$ comprise the ones of type II. Concretely, a basic classical Lie superalgebra of type I is isomorphic to one of the following Lie superalgebras: $A(m,n)=\mathfrak{sl}(m+1|n+1)^{\mathbb{C}}$ with $m>n\geq 0$; $A(n|n)=\mathfrak{psl}(n+1|n+1)^{\mathbb{C}}$ with $n\geq 1$; or $C(n)=\mathfrak{osp}(2|2n-2)^{\mathbb{C}}$ with $n\geq 3$ (noting that $C(2)\cong A(1|0)$).
A real Lie superalgebra $\mathfrak{g}$ is called a real form of a complex Lie superalgebra $\mathfrak{g}^{\mathbb{C}}$ if $\mathfrak{g}^{\mathbb{C}}\cong\mathfrak{g}\otimes_{\mathbb{R}}\mathbb{C}$, while a star operation $\mathfrak{g}^{\mathbb{C}}\to\mathfrak{g}^{\mathbb{C}}$ is an even anti-linear map such that
@@
for any $X,Y\in\mathfrak{g}^{\mathbb{C}}$; see [SNR77] or [DM99, §4.1–4.3]. Here, anti-linearity means that $(cX)^{*}=\bar{c}X^{*}$, where $\bar{c}$ is the complex conjugate of $c\in\mathbb{C}$. With $\sqrt{\imath}=e^{\frac{\pi}{4}\imath}$, a real form $\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}}$ can now be constructed as
$\begin{gathered}\mathfrak{g}_{\bar{0}}=\operatorname{span}_{\mathbb{R}}\{X\in\mathfrak{g}_{\bar{0}}^{\mathbb{C}}\,|\,X^{*}=-X\},\\[2.84544pt] \mathfrak{g}_{\bar{1}}=\operatorname{span}_{\mathbb{R}}\{\sqrt{\imath}X\in\mathfrak{g}_{\bar{1}}^{\mathbb{C}}\,|\,X^{*}=-X\}=\operatorname{span}_{\mathbb{R}}\{X\in\mathfrak{g}_{\bar{1}}^{\mathbb{C}}\,|\,X^{*}=\imath X\}.\end{gathered}$ (2.4)
Note that $\mathfrak{g}_{\bar{0}}$ is a real form of the complex Lie algebra $\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$. We say that a real form of $\mathfrak{g}^{\mathbb{C}}$ is compact if $\mathfrak{g}_{\bar{0}}$ is a compact Lie algebra; cf. [AN15, FG23]. It was shown in [SNR77] that among basic classical Lie superalgebras, only type I Lie superalgebras admit star operations. In that case, there are actually two star operations, and they are related by duality. Throughout this paper, we are only concerned with the so-called type (1) star operation; see, e.g., [GZ90a, GZ90b] and Section 4.2.
@@ 2.2 Lie supergroups
for any open set $U\subseteq\mathbb{R}^{m}$. Here, $\mathcal{C}_{\mathbb{R}^{m}}^{\infty}$ denotes the sheaf of smooth (even) functions on $\mathbb{R}^{m}$, and $\Lambda_{\mathbb{R}}[\xi_{1},\dots,\xi_{n}]$ is the Grassman algebra in the odd variables $\xi_{1},\dots,\xi_{n}$.
A (real smooth) supermanifold of dimension $m|n$ is a ringed space $M=(|M|,\mathcal{O}_{M})$, where the underlying topological space $|M|$ is Hausdorff and second-countable, and $\mathcal{O}_{M}=(\mathcal{O}_{M})_{\bar{0}}\oplus(\mathcal{O}_{M})_{\bar{1}}$ is a sheaf of superfunctions, i.e., for each $x\in|M|$ there exists an open neighbourhood $V_{x}\subseteq|M|$ of $x$ and an open set $U\subseteq\mathbb{R}^{m|n}$ such that $\mathcal{O}_{M}(V_{x})\cong\mathcal{O}_{\mathbb{R}^{m|n}}(U)$. Let $\mathcal{I}_{M}=\langle(\mathcal{O}_{M})_{\bar{1}}\rangle$ be the two-sided ideal of $\mathcal{O}_{M}$ generated by all nilpotent sections in $(\mathcal{O}_{M})_{\bar{1}}$. The corresponding subspace $M_{0}=(|M|,\mathcal{O}_{M}/\mathcal{I}_{M})$ is called the reduced manifold associated with $M$, where $\mathcal{O}_{M}/\mathcal{I}_{M}\cong\mathcal{C}^{\infty}_{|M|}$ is regarded as the sheaf of smooth functions on $|M|$. Hence, $M_{0}$ lies in the category of ordinary smooth manifolds. The canonical projection $\mathrm{ev}:\mathcal{O}_{M}\to\mathcal{O}_{M}/\mathcal{I}_{M}$ is the corresponding evaluation map. For any point $p\in|M|$, let $\mathcal{O}_{M,p}$ denote the stalk at $p$. The evaluation at $p$ is then defined as
$\mathrm{ev}_{p}:\mathcal{O}_{M,p}\to\mathbb{R},\qquad f\mapsto f(p):=\mathrm{ev}(f)(p).$
@@
$\varphi(fg)=\varphi(f)g+(-1)^{[\varphi][f]}f\varphi(g),\qquad f,g\in\mathcal{O}_{M}(U).$
The sections of $\mathcal{T}_{M}$ are called vector fields. For every point $p\in|M|$, the tangent superspace $T_{p}M$ at $p$ is the vector superspace of dimension $m|n$, defined by
$T_{p}M:=\{\psi\in\mathrm{Hom}_{\mathbb{R}}(\mathcal{O}_{M,p},\mathbb{R})\,|\,\psi(fg)=\psi(f)g(p)+(-1)^{[\psi][f]}f(p)\psi(g)\}.$
Elements of $T_{p}M$ are called tangent vectors. For any $p\in U$ and vector field $X\in\mathcal{T}_{M}(U)$, the value of $X$ at $p$ is the tangent vector $X_{p}:\mathcal{O}_{M,p}\to\mathbb{R}$ defined by $X_{p}(f):=\mathrm{ev}_{p}(X(f))$ for all $f\in\mathcal{O}_{M,p}$. A vector field is only determined by its values at all points if the odd dimension of $M$ is zero. We also introduce the dual space
$\mathcal{T}_{M}^{*}:=\mathrm{Hom}_{\mathcal{O}_{M}}(\mathcal{T}_{M},\mathcal{O}_{M}).$
@@
$v\mapsto X_{v}=(1\otimes v)\circ m^{*},$
where $X_{v}$ is the left invariant vector field with $(X_{v})_{e}=v$. We refer to [CCF11] for more details.